David has a total of 4.10 pesos in 5 cent, 10 cent and 25 cent coins. He has twice as many 5 cent coins and two more 25 cent coins than 10 cent coins. How many 10 cent coins does he have?
Answer by ikleyn(52788)

.

Coins:  10-cent   x

         5-cent   2x

        25-cent   x+2


Money equation


    10x + 5*(2x) + 25*(x+2) = 410  cents


    10x + 10x + 25x + 50 = 410

    45x                  = 410-50 = 360

      x                           = 360/45 = 8


ANSWER.  8  of the 10-cent coins;  8*2 = 16  of the 5 cent-coins  and  8+2 = 10 25-cent coins.


CHECK.   8*10 + 16*5 + 25*10 = 410 cents.   ! Correct !

Solved.

---------------------

From my post, learn how to solve such problems using one single variable from the very beginning to the end.